OpenStudy (august899):

Which angle is not coterminal to 4pi/3 radians? -120 degrees,960 degrees,-540 degrees,or -840 degrees?

OpenStudy (jdoe0001):

well, convert the radian unit to degrees, and then check which one isn't coterminal to it :) \(\begin{array}{llll} degrees&radians \\\hline\\ 180&\pi \\ x&\frac{4\pi }{3} \end{array}\implies \cfrac{180}{x}=\cfrac{\pi }{\frac{4\pi }{3}}\)